# Chemical Language Model Linker: blending text and molecules with modular adapters

### v1

@@ 5 Discussion
This study proposes the ChemLML framework for using pretrained language and molecular models in text-guided molecule design. By reusing large-scale pretrained models, we enhance the flexibility of text-based molecule generation and reduce the training effort by supporting multiple types of frozen text encoders. In our case study generating candidate inhibitors for three protein targets from text, we find that for the two targets where docking performs well, ChemLML T5 encoder+MolGen generates molecules with docking scores that are better than ChemLML T5 encoder+MolXPT, two existing molecule generation algorithms, and two distributions of control molecules. For the inosine-5’-monophosphate dehydrogenase target, the median ChemLML docking score is even better than the ground truth inhibitor from PubChem.
When finetuning LLMs, we find it is easy to finetune Galactica 125M. However, it becomes harder to finetune the Galactica 1.3B model. Other finetuning methods such as Low-Rank Adaptation (Hu et al., 2022) may solve this problem. Also, we do not carry out experiments on Galactica 30B and 120B due to hardware and training technique limitations. Improving the LLM finetuning and finetuning larger LLMs are directions for future work.
A limitation of ChemLML is that it only focuses on molecule generation whereas related methods are multi-task. For instance, MolT5 can perform molecule captioning, and Text+ChemT5 can also conduct forward reaction and retrosynthesis prediction. Conceptually, the ChemLML framework is compatible with a multi-task setting and even more modalities including proteins (Liu et al., 2023b, 2024). This presents another avenue for future work.
@@
Many molecule generation models, including ChemLML, still face limitations in generating fully plausible, semantically-valid molecules even when then are syntactically-valid. We observed that ChemLML-generated molecules in the SMILES or SELFIES format may fail to be transformed to 3D conformers for docking. In general, inspection of other molecule generation methods has uncovered unstable ring systems, reactive or toxic moieties, or geometric errors (Walters, 2024). ChemLML may be susceptible to these issues as well.
Finally, generating structurally similar molecules does not necessarily ensure that they will share the same properties as the target molecule. Many molecule descriptions pertain to the chemical structure of the molecules. In these cases, the assumption made by the chemical structure-based evaluations is most appropriate. For other physicochemical properties, different chemical structures could yield similar properties. In these cases, the selection of improved evaluated metrics remains another area for future exploration.
@@ 6 Conclusion
ChemLML introduces the strategy of using an adapter between text LLMs and molecule decoders for text-guided molecule design. Our approach is designed to capitalize on the rampant advances in both natural language modeling and unconditional molecule generation. The adapter is lightweight and compatible with different kinds of pretrained LLM encoders and molecule decoders. Looking forward, the biological and chemical communities have shown strong interest in using natural language to condition the generation of small molecules and even proteins with desired structural and functional properties Pei et al. (2024). ChemLML provides a path for combining the most advanced natural language and domain-specific models as they continue to undergo rapid development for this type of text-conditioned generation.
Acknowledgments
This research was supported by National Institutes of Health award R01GM135631. The research was performed using the computing resources and assistance of the University of Wisconsin-Madison Center for High Throughput Computing (Center for High Throughput Computing, 2006). We thank Shengchao Liu for addressing problems with dataset collection, Yin Fang for discussion of Mol-Instruction, and Chaowei Xiao for insightful discussion about the experimental setting.

### v2

@@ 5 Discussion
This study proposes the ChemLML framework for combining pretrained natural language and molecular models in text-guided molecule design. By reusing large-scale pretrained models, we enhance the flexibility of text-based molecule generation and reduce the training effort by supporting multiple types of frozen text encoders. In our case study generating candidate inhibitors for eight protein targets from text, we find that ChemLML T5 encoder+MolGen often generates molecules with docking scores that are better than ChemLML T5 encoder+MolXPT, two existing molecule generation algorithms, and two distributions of control molecules. For four targets, the median ChemLML docking score is even better than the ground truth inhibitor from PubChem.
When finetuning LLMs, we find it is easy to finetune Galactica 125M. However, it becomes harder to finetune the Galactica 1.3B model. Other finetuning methods such as Low-Rank Adaptation 67 may solve this problem. Also, we do not carry out experiments on Galactica 30B and 120B due to hardware and training technique limitations. Improving the LLM finetuning and finetuning larger LLMs are directions for future work.
A limitation of ChemLML is that it only focuses on molecule generation whereas related methods are multi-task. For instance, MolT5 can perform molecule captioning, and Text+ChemT5 can also conduct forward reaction and retrosynthesis prediction. Mol-Instruction 41 introduced multi-task dataset that supports a wide range of molecular learning tasks. Conceptually, the ChemLML framework is compatible with the multi-task setting and even more modalities including proteins 68, 51. This presents another avenue for future work.
@@
Many molecule generation models, including ChemLML, still face limitations in generating fully plausible, semantically-valid molecules even when then are syntactically-valid. We observed that ChemLML-generated molecules in the SMILES or SELFIES format may fail to be transformed to 3D conformers for docking. In general, inspection of other molecule generation methods has uncovered unstable ring systems, reactive or toxic moieties, or geometric errors 69. ChemLML may be susceptible to these issues as well. Alternative molecular representations beyond SMILES and SELFIES 70 could be relevant for addressing these issues in the future as well as additional careful assessments of the impact of the choice of molecular representation 60.
Finally, generating structurally similar molecules does not necessarily ensure that they will share the same properties as the target molecule. Many molecule descriptions pertain to the chemical structure, i.e., bond topology or scaffold geometry, of the molecules. In these cases, the assumption made by the chemical structure-based evaluations, like fingerprint similarity or docking, is most appropriate. For other physicochemical properties, like water solubility, different chemical structures could yield similar properties. In these cases, the selection of improved evaluated metrics remains another area for future exploration.
@@ 6 Conclusion
ChemLML introduces the strategy of using an adapter between text LLMs and molecule decoders for text-guided molecule design. Our approach is designed to capitalize on the rampant advances in both natural language modeling and unconditional molecule generation. The adapter is lightweight and compatible with different kinds of pretrained LLM encoders and molecule decoders. Looking forward, the biological and chemical communities have shown strong interest in using natural language to condition the generation of small molecules and even proteins with desired structural and functional properties 45. ChemLML provides a path for combining the most advanced natural language and domain-specific models as they continue to undergo rapid development for this type of text-conditioned generation.
7 Acknowledgments
This research was supported by National Institutes of Health award R01GM135631. The research was performed using the computing resources and assistance of the University of Wisconsin-Madison Center for High Throughput Computing 71. We thank Shengchao Liu for addressing problems with dataset collection, Yin Fang for discussion of Mol-Instruction, and Chaowei Xiao for insightful discussion about the experimental setting.